OpenStudy (anonymous):

Jack's age next year will be twice Jill's age last year. Their present ages total 45. How old is each now?

OpenStudy (anonymous):

Jack's age is \(y\) and Jill's age is \(x\). "Jack's age next year will be twice Jill's age last year" How do we write this using our variables?

OpenStudy (anonymous):

Here is a hint: "Jack's age next year" \(y+1\) "Jill's age last year" \(x-1\) So: "\(y+1\) will be twice \(x-1\)"

OpenStudy (anonymous):

You need to start turning this into an equation first.

OpenStudy (anonymous):

We don't want any words to remain.

OpenStudy (anonymous):

y-1=2(x-1)

OpenStudy (anonymous):

Good. Now how do you write: "Their present ages total 45."

OpenStudy (anonymous):

Hint: total means we need to do addition.

OpenStudy (anonymous):

y-1=2(x-1)+45

OpenStudy (anonymous):

Nope, this is a separate equation. We don't need to worry about the first equation right now.

OpenStudy (anonymous):

"Their present ages total 45." Is a separate equation. Hint: it will have \(y+x\) in it.